OpenStudy (anonymous):

What is the logarithmic form of the equation e4x ≈ 2981? log4x2981 = e 4 logxe = 2981 ln 4x = 2981 ln 2981 = 4x

OpenStudy (anonymous):

@jhonyy9 i need help again

jhonyy9 (jhonyy9):

there is e^(4x) ?

OpenStudy (anonymous):

is it going to be In(2981) 4x

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

e^4x=2981

jhonyy9 (jhonyy9):

do you know what is the base of ln

OpenStudy (anonymous):

yep i got the answer thank u again...might need so more help

jhonyy9 (jhonyy9):

so what will be right sure and how you get it ?

OpenStudy (anonymous):

In(2981) 4x

jhonyy9 (jhonyy9):

ln has the e like base - hope so much that you know - yes ?

jhonyy9 (jhonyy9):

so there are e^(4x) = 2981 make logarithm naturale on the both sides and what will get ?

jhonyy9 (jhonyy9):

ln e^(4x) = ln 2981 yes ? can you continue it ?

jhonyy9 (jhonyy9):

so how i have said above the logarithm natural has the base e what mean that ln e =1 so in this case there are ln e^(4x) = ln 2981 4xln e = ln 2981 where you know that ln e = 1 than will result 4x = ln 2981 hope so much that is understandably sure right easy yes ?

jhonyy9 (jhonyy9):

the last one from these choices right ?

jhonyy9 (jhonyy9):

sorry just inverse - ok. ?

jhonyy9 (jhonyy9):

@Help100